OpenStudy (anonymous):

Let A be an 80  80 matrix. Prove that if AA^T = I, then det(A) 1 = or -1.

OpenStudy (anonymous):

.

OpenStudy (zarkon):

is that 80x80 matrix ??

OpenStudy (anonymous):

det(AB) = det(A)det(B) det(A^T) = det(A) put this all together : det(AA^T) = det(A)det(A^T) = det(A)det(A) = det(I) det(A)^2 = 1 det(A) = 1,-1

OpenStudy (anonymous):

yes, 80x80

OpenStudy (zarkon):

Coolsector's method works for any nxn

OpenStudy (anonymous):

zarkon...i posted question for u again